OpenStudy (anonymous):

Help!!! how do you simplify a problem like this (5/3)^-2????

OpenStudy (callisto):

\[x^{-a} = 1/(x^a)\](5/3)^-2 = 1/ [(5/3)^2] = 1 / (25/9) Can you do it from here?

OpenStudy (anonymous):

umm i think so thankz

OpenStudy (callisto):

welcome :)